Large Consignments of computer components are inspected for
defectives by means of a sampling system. Ten components are
examined and the lot is to be rejected if two or more are found
to be defective. If a consignments contains exactly 10%
defectives.
Find the probability of the consignment by using the technique
of Binomial probability distribution that the consignment is:
i) Accepted
ii) Rejected
1
Expert's answer
2021-07-19T07:32:15-0400

Binomial (10,0.1)

"P(x)=\binom{n}{x}P^x(1-P)^{n-x}, x=0,1,2,...,10"

i. Probability of accepting is "P(x\le1)"

"P(x\le1)=P(x=0)+P(x=1)"

"=[\binom{10}{0}\times(1-0.1)^{10}]+[\binom{10}{1}\times0.1^1(1-0.1)^9]"

"=0.349+0.387"

"=0.736"

ii. Probability of rejecting is "P(x>1)"

"P(x>1)" is 1- probability of accepting

"P(x>1)=1-P(x\le1)"

"=1-0.736"

"=0.264"